OpenStudy (anonymous):

Given y inversely proportional to x and x = 3 for y = 4, what is x if y = 12?

OpenStudy (anonymous):

y=k*(1/x) 4=k(1/3)

OpenStudy (jdoe0001):

"y" is inverse proportional to "x" the inverse of "x" is 1/x

OpenStudy (anonymous):

solve for k then 12=k*(1/x) I think.

OpenStudy (dan815):